# Algorithmic Aspects of Strategic Trading

### v1

@@ 6 Experiments
In light of our theoretical results, it is natural to ask how quickly no-regret dynamics converge to an approximate coarse correlated equilibrium in actual implementation, and what the approximate equilibria look like — in particular, are they “close” to the stronger notion of Nash equilibria? We empirically investigate these questions under different regimes of market impact.
@@
We implement no-regret dynamics between two players using our instantiation of FTPL (Algorithm 3) as described in Section 5. Throughout, we will fix the setting of $T=5,V_{1}=V_{2}=10,\theta_{L}=-5,\theta_{U}=5$ (thus both players have 5 time periods in which to acquire a net long position of 10 shares, and are able to buy or (short) sell up to 5 shares at each step) while varying the market impact coefficient $\kappa$. For each setting of $\kappa$ we execute 100 runs of no-regret dynamics, each consisting of 2500 rounds. Recall that FTPL takes in an additional parameter: the noise parameter $\eta$. Using the theoretical guideline of $\eta\approx\sqrt{\text{number of rounds}}$, we choose $\eta=50$.
@@ 6.1 Convergence Rate
Figure 4: Average/per-round regrets of players 1 and 2 for varying $\kappa$, averaged across 100 runs of no-regret dynamics. We exclude other $\kappa$ for conciseness; the curves (as shown in this manner) look fairly identical.
@@ 6.2 Equilibria Properties
We have established that no-regret dynamics converge to approximate coarse correlated equilibria fairly quickly. Now we ask: What do the approximate equilibria found by no-regret dynamics look like? Can we say anything more interesting or specific about these equilibria; in particular, are they close to Nash equilibria?
Recall that a coarse correlated equilibrium is a mixed Nash equilibria if its joint distribution can be written as a product distribution—that is, each player’s actions can be drawn independently from their own marginal distributions. Since the strategy spaces are not numeric but discrete, combinatorial objects, we cannot measure correlations between player actions in the standard way, but instead will examine the total variation (TV) distance between the joint distribution returned by no-regret dynamics and the product of each player’s marginal distribution. More specifically, let $\mathbf{D}$ be the empirical joint distribution over the realized action pairs $(a_{1,1},a_{2,1}),...,(a_{1,R},a_{2,R})$. Let $D_{1}$ be the marginal distribution over the first player’s actions and $D_{2}$ be the marginal distribution over the second player’s actions. We compute the TV distance between $\mathbf{D}$ and $D_{1}\times D_{2}$ as follows:
@@
In the sequel, we will refer to this distance informally as “correlation” between player strategies. Table 1 shows TV distances for varying $\kappa$. For each setting of $\kappa$, we report the average TV distance computed over 100 runs of no-regret dynamics. As expected, TV distance is low for the special case of $\kappa=2$ (no-regret dynamics are known to converge to Nash equilibria in zero/constant-sum games). The TV distance is particularly low for $\kappa=0$ — in our next set of results, we see that this can be explained by the fact that when $\kappa=0$, no-regret dynamics finds a pure Nash equilibrium fairly quickly (recall that pure Nash equilibria are guaranteed to exist in this regime). For large $\kappa$, the approximate coarse correlated equilibria found by no-regret dynamics exhibit high correlation.
$\kappa$ 0 0.5 1 1.5 2 2.5 3 5 10
mean 0.033 0.211 0.216 0.475 0.181 0.971 1.028 0.958 0.728
standard deviation 0.003 0.028 0.016 0.014 0.054 0.039 0.062 0.097 0.168
Table 1: TV distances between outputted joint distribution and product of marginal distributions, for varying $\kappa$. The table shows means and standard deviations over 100 runs of no-regret dynamics for each $\kappa$.
Figure 5: Welfare of approximate CCE for varying $\kappa$. For each $\kappa$, we plot the average welfare computed over 100 runs of no-regret dynamics.
@@
Finally, we touch on the social welfare of the equilibria found by no-regret dynamics. Recall that expected welfare is defined as $\sum_{i=1}^{n}\mathbb{E}_{\bm{a}\in\mathbf{D}}[c(a_{i},a_{-i})]$ (since we are talking about players’ costs, lower welfare is better). Figure 5 shows the welfare of the approximate coarse correlated equilibria found by no-regret dynamics for varying $\kappa$. In general, we would expect costs — and thus welfare — to increase as $\kappa$ increases. Interestingly, we see an inflection point at $\kappa=2$; welfare increases at a slower rate for large $\kappa$. Previously, we observed that for large $\kappa$, no-regret dynamics returns equilibria exhibiting high correlation (i.e. high TV distances). We can interpret this as preliminary evidence showing that allowing for correlation might improve social welfare in markets where permanent impact dominates.

### v2

@@ 6 Experiments
In light of our theoretical results, it is natural to ask how quickly no-regret dynamics converge to an approximate CCE in actual implementation, and what the approximate equilibria look like — in particular, are they “close” to the stronger notion of Nash equilibria? We empirically investigate these questions under different regimes of market impact, which dictate how “close” or “far’” the game is from a potential game and a zero-sum game. The code can be found here.
@@
We implement no-regret dynamics between two players using our instantiation of FTPL as described in Section 5. Throughout, we will fix the setting of $T=5,V_{1}=V_{2}=10,\theta_{L}=-5,\theta_{U}=5$ (thus both players have 5 time periods in which to acquire a net long position of 10 shares, and are able to buy or (short) sell up to 5 shares at each step) while varying the market impact coefficient $\kappa$. For each setting of $\kappa$ we execute 100 runs of no-regret dynamics, each consisting of 2500 rounds. Recall that FTPL takes in an additional noise parameter $\eta$. Using the theoretical guideline of $\eta\approx\sqrt{\text{number of rounds}}$, we choose $\eta=50$.
@@ 6.1 Convergence Rate
Figure 4: Average/per-round regrets of players 1 and 2 for varying $\kappa$, averaged across 100 runs of no-regret dynamics. We exclude other $\kappa$ for concision; the curves (as shown in this manner) look fairly identical.
@@ 6.2 Equilibria Properties
We have established that no-regret dynamics converge to approximate coarse correlated equilibria fairly quickly. Now we ask: What do the approximate equilibria found by no-regret dynamics look like? Can we say anything more interesting or specific about these equilibria; in particular, are they close to stronger forms of equilibria like Nash equilibria?
Figure 5 answers the questions: how close is the outputted joint distribution to the stronger notion of CE, and how close are the outputted marginal distributions to a mixed Nash equilibrium? We measure these distances directly using the definitions:
Dist. to NE $=\max_{i=1,2}\left[\mathbb{E}_{\begin{subarray}{c}a_{1}\sim D_{1}\\ a_{2}\sim D_{2}\end{subarray}}[c(a_{i},a_{-i})]-\min_{a^{*}\in\mathcal{A}}\mathbb{E}_{a_{-i}\sim D_{-i}}[c(a^{*},a_{-i})]\right]$
Dist. to CE $=\max_{i=1,2}\left[\mathbb{E}_{(a_{1},a_{2})\sim\mathbf{D}}[c(a_{i},a_{-i})]-\min_{\phi:\mathcal{A}\to\mathcal{A}}\mathbb{E}_{(a_{1},a_{2})\sim\mathbf{D}}[c(\phi(a_{i}),a_{-i})]\right]$
Figure 5: Distances to Nash equilibria and CE for varying $\kappa$. We use distance to CCE (regret) as a baseline. The plots show means and std. deviations over 100 runs of no-regret dynamics. We also plot the distance to CE found by no-swap-regret dynamics (means and std. deviations over 20 runs of a no-swap-regret algorithm; see Appendix A).
Our results show that when the game interpolates between a potential game and zero-sum game ($\kappa\leq 2$), FTPL finds almost-exact CE and mixed Nash equilibria; in fact the distances closely mirror the distances to CCE, which is theoretically guaranteed to be low. When the game is zero-sum (i.e. when $\kappa=2$), no-regret dynamics is guaranteed to converge to a mixed Nash equilibrium (Cesa-Bianchi and Lugosi, 2006). Thus for $\kappa\geq 2$, the figure presents an intuitive relationship: as the game becomes less like a zero-sum game (i.e. as $\kappa$ increases beyond $2$), the distance to NE found by FTPL dynamics also increases. Meanwhile, for larger $\kappa$, the distance to CE increases sharply then plateaus.
Figure 5 also shows the distance to to CE of the joint distribution outputted by a recent no-swap-regret algorithm (Dagan et al., 2024; Peng and Rubinstein, 2024), which guarantees convergence to an approximate CE. We see that although FTPL dynamics only guarantees convergence to approximate CCE, its joint play is in fact closer to CE than an implementation of no-swap-regret dynamics for all $\kappa$. We formally present the no-swap-regret algorithm and give implementation details in Appendix A.
Figure 6: TV distances between outputted joint distribution and product of marginal distributions, for varying $\kappa$. The plot shows means and std. deviations over 100 runs of no-regret dynamics.
Next we measure how “correlated” the joint history of play is. Recall that a CCE is a mixed Nash equilibria if its joint distribution can be written as a product distribution—that is, each player’s actions can be drawn independently from their own marginal distributions. Since the strategy spaces are not numeric but discrete, combinatorial objects, we cannot measure correlations between player actions in the standard way, but instead will examine the total variation (TV) distance between the joint distribution returned by no-regret dynamics and the product of each player’s marginal distribution. More specifically, let $\mathbf{D}$ be the empirical joint distribution over the realized action pairs $(a_{1,1},a_{2,1}),...,(a_{1,R},a_{2,R})$. Let $D_{1}$ be the marginal distribution over the first player’s actions and $D_{2}$ be the marginal distribution over the second player’s actions. We compute the TV distance between $\mathbf{D}$ and $D_{1}\times D_{2}$ as follows:
@@
In the sequel, we will refer to this distance informally as “correlation” between player strategies. Figure 6 shows TV distances for varying $\kappa$. For each setting of $\kappa$, we report the average TV distance computed over 100 runs of no-regret dynamics. As expected, TV distance is low for the special case of $\kappa=2$ (no-regret dynamics are known to converge to Nash equilibria in zero/constant-sum games). The TV distance is particularly low for $\kappa=0$ — in our subsequent results, we see that this can be explained by the fact that when $\kappa=0$, no-regret dynamics finds a pure Nash equilibrium fairly quickly (recall that pure Nash equilibria are guaranteed to exist in this regime). For large $\kappa$, the approximate coarse correlated equilibria found by no-regret dynamics exhibit high correlation.
@@
Figure 8: Welfare of approximate CCE for varying $\kappa$. For each $\kappa$, we plot the mean welfare computed over 100 runs of no-regret dynamics (std deviation bars are too small to appear on plot).
Finally, we touch on the social welfare of the equilibria found by no-regret dynamics. Recall that expected welfare is defined as $\sum_{i=1}^{n}\mathbb{E}_{\bm{a}\in\mathbf{D}}[c(a_{i},a_{-i})]$ (since we are talking about costs, lower welfare is better). Figure 8 shows the welfare of the approximate coarse correlated equilibria found by no-regret dynamics for varying $\kappa$. In general, we would expect costs — and thus welfare — to increase as $\kappa$ increases. Interestingly, we see an inflection point at $\kappa=2$; welfare increases at a slower rate for large $\kappa$. Previously, we observed that for large $\kappa$, no-regret dynamics returns equilibria exhibiting high correlation (i.e. high TV distances). We can thus interpret this as preliminary evidence showing that allowing for correlation might improve social welfare in markets where permanent impact dominates. Indeed, we can think of correlation as a form of collusion. Our findings suggest that as $\kappa$ increases—intuitively, this is also as we move farther from a zero-sum game—the potential benefits of collusion increase. Moreover, this type of collusive behavior can emerge organically from simple learning dynamics.
7 Limitations and Future Work
Our work leaves open several avenues for future work. In this work, we considered static strategies. Although we can view our static strategies as baselines or planned trajectories for more sophisticated strategies, future research could explore competitive trading with dynamic strategies that adapt to market activity. Other interesting directions include more general cost functions, such as non-linear functions and those modeling decaying permanent impact, and incorporating different types of (possibly unknown) players with objectives beyond pure position acquisition. Finally, in our game, the parameter $\kappa$ governs how close the game is to a potential game and a zero-sum game. Given that every game can be decomposed into a potential and a zero-sum game, it is broadly relevant to investigate how equilibria properties and computation evolve with shifting game structure, independent of our specific trading context.
@@ 8 Acknowledgements
Appendix A Correlated Equilibria Computation
In Section 5, we gave an efficient algorithm to compute approximate coarse correlated equilibria via no-regret dynamics. Here, we investigate efficient computation of correlated equilibria (CE), a stronger equilibrium concept than CCE.
Just as coarse correlated equilibrium is tied to the notion of regret (otherwise referred to as external regret), correlated equilibrium is tied to the notion of swap regret.
Definition A.1 (Swap Regret).
Let $\Phi_{i}=\{\phi:\mathcal{A}_{i}\to\mathcal{A}_{i}\}$ be the collection of all functions mapping actions to actions. The (average) swap regret of a player $i$ who chooses a sequence of actions $a_{i,1},...,a_{i,R}\in\mathcal{A}_{i}$ is defined as:
$SReg_{i}(R)=\max_{\phi\in\Phi_{i}}\frac{1}{R}\sum_{r=1}^{R}(c(a_{i,r},a_{-i,r})-c(\phi(a_{i,r}),a_{-i,r}))$
No swap regret is a stronger guarantee than no (external) regret; it asks that given the history of play, a player has no incentive to deviate to a fixed strategy conditioned on the strategy they chose. By definition, if every player at swap regret at most $\varepsilon$, then the empirical joint distribution of play is an $\varepsilon$-approximate correlated equilibrium.
To implement no-swap-regret dynamics, we use an existing reduction transforming any no-regret algorithm to a no-swap-regret algorithm. The classical reduction of Blum and Mansour (Blum and Mansour, 2007) guarantees, against any sequence of opponent actions, $SReg(R)\leq|\mathcal{A}|\cdot Reg(R)$ given an algorithm that obtains regret bounded by $Reg(R)$ (for certain no-regret algorithms, a tighter analysis improves the dependence on $|\mathcal{A}|$ by a factor of $\sqrt{|\mathcal{A}|}$). To handle a large action space, we use recent reductions of Dagan et al. (2024) and Peng and Rubinstein (2024) that guarantee vanishing swap regret at a rate depending only on the external regret guarantee of the no-regret algorithm, avoiding any dependence on the number of actions (however this comes at an exponential cost in the approximation parameter). We state its guarantees below but defer details of the reduction to Dagan et al. (2024) and Peng and Rubinstein (2024).
Theorem A.2 (Theorem 3.1 of Dagan et al. (2024)).
Fix an action set $\mathcal{A}$. Fix $M,d,R\in\mathbb{N}$ such that $M^{d-1}\leq R\leq M^{d}$. Given an algorithm that, against any adversarial sequence of actions, guarantees regret at most $Reg(R)$ after $R$ rounds, there is an algorithm producing randomized actions $p_{1},...,p_{R}\in\Delta\mathcal{A}$ such that (in expectation over the randomized actions):
$SReg(R)\leq Reg(M)+\frac{3}{d}$
Moreover, if the per-round running time of the no-regret algorithm is $C$, then the per-round amortized running time of this algorithm is $O(C)$.
Importing our instantiation of FTPL, we obtain the following guarantee on the joint history given by no-swap-regret dynamics. The number of rounds needed to reach an $\varepsilon$-approximate correlated equilibrium is polynomial in the parameters of the game, but depends exponentially on $1/\varepsilon$.
Corollary A.3.
Fix a player $i$ in the trading game with action set $\mathcal{A}_{i}$. Using the instantiation of FTPL given by Corollary 5.3, there is an algorithm producing randomized actions $p_{1},...,p_{R}\in\Delta\mathcal{A}_{i}$ such that, in expectation over the randomized actions, $SReg_{i}(R)\leq\varepsilon$ after $R=O\left(\left(\frac{n^{2}\theta^{5}T^{6}}{\varepsilon^{2}}\right)^{\frac{1}{\varepsilon}}\right)$ rounds. Moreover, the per-round amortized running time is $O(\theta^{2}T^{2})$, where $\theta=\max\{|\theta_{L}|,|\theta_{U}|\}$. If every player $i$ runs a copy of this algorithm, the empirical distribution $\mathbf{D}$ is an $\varepsilon$-approximate correlated equilibrium after $R=O\left(\left(\frac{n^{2}\theta^{5}T^{6}}{\varepsilon^{2}}\right)^{\frac{1}{\varepsilon}}\right)$ rounds of no-swap-regret dynamics.
Proof.
We plug the regret guarantee of FTPL from Corollary 5.3, $M=\frac{n^{2}\theta^{5}T^{6}}{\varepsilon^{2}}$, and $d=\frac{1}{\varepsilon}$ into Theorem A.2. The runtime complexity follows from Corollary 5.4 and Theorem A.2. ∎
A.1 Details of Experimental Implementation
We implement the no-swap-regret algorithm of Corollary A.3 and plot the distance to CE (i.e. swap regret) of the joint play in Figure 5. In our implementation, we use the same game parameters as before: $T=5,V_{1}=10,V_{2}=10,\theta_{L}=-5,\theta_{U}=5$. Given that $M\gg d$ in theory, we set $M=150$ and $d=2$. For each setting of $\kappa$, we execute 20 runs of no-swap-regret-dynamics, each consisting of $M^{d}=22,500$ rounds.
The results suggest that in practice, the algorithm is indeed hindered by the (exponentially) slow convergence rate given by theory. In fact, we find that FTPL obtains lower swap regret than this algorithm (it only guarantees low external regret), even with much fewer rounds.